OpenStudy (anonymous):

Please help me solve this its all factoring. medal will be given thanks. 1. 2y^2-3^z 2. 4x^2-y^2+2y-1 3. 16p^2-(5q-r)^2 4. 4x^3+40x^2+100x 5. m^6-n^6

OpenStudy (jack1):

Q2. (2x - y - 1) (2x + y - 1)

OpenStudy (jack1):

Q1. I dont think you can factor this any further...

OpenStudy (anonymous):

thanks jack can you put the solution also it would help me a lot so i can review it.

OpenStudy (unklerhaukus):

For question 4. \[4x^3+40x^2+100x\] you can factor out \(4x\) from each term

OpenStudy (jack1):

y = ax^2 + by^2 + cy + d so 4x^2-y^2+2y-1 = 4x^2 - (y-1)^2 = (2x - y - 1) (2x + y - 1)

OpenStudy (anonymous):

thank you UnkleRhaukus i will try to solve it and Jack1 appreciated.

OpenStudy (jack1):

Q3 16p^2-(5q-r)^2 the 16p^2 = 4p x 4p and the (5q-r) ^2 = (5q-r) (5q+r) so = (4p +5q - r) (4p -5q + r)

OpenStudy (jack1):

Q5... I'm out...? sorry i know it will be a large combination of (m - n) (m + n)...?

OpenStudy (unklerhaukus):

for 5. \[m^6-n^6\] it helps to temporarily set \[m^3=M, \qquad n^3=N\]

OpenStudy (unklerhaukus):

so factorise \(M^2-N^2\) = ...

OpenStudy (unklerhaukus):

for 1. im not sure if you typed it correctly?

OpenStudy (anonymous):

thanks you unklerhaukus 1. is written mistakely its 2y^2-3z im sorry for that.

OpenStudy (unklerhaukus):

hmm, well 1. still can not be factored.

OpenStudy (anonymous):

thanks you unkle im currently answering the other no ill post later with the answers